OpenStudy (anonymous):

Use Euler's method with step size 0.2 to estimate y(1.4), where y(x) is the solution of the initial-value problem y'=x-xy, y(1)=0

OpenStudy (anonymous):

\( y'=f(x,y)=x-xy \) and the formula is \( y_{n+1}=y_{n}+h f(x_{n} , y_{n}) \)

OpenStudy (anonymous):

oh ok. Thanks

OpenStudy (anonymous):

h is your step size =0.2 \( x_{0}=1 \ \ y_{0}=0 \)

OpenStudy (anonymous):

What's the big picture? What am I trying to do here?

OpenStudy (anonymous):

well thats a numerical method for solving differential equations and coming up from taylor series http://en.wikipedia.org/wiki/Numerical_methods_for_ordinary_differential_equations

OpenStudy (anonymous):

\[y_1=y_0+0.2f(1,0)\]

OpenStudy (anonymous):

what's my \[y_0\]?

OpenStudy (anonymous):

y0=0

OpenStudy (anonymous):

\[y_1=0+0.2f(1,0)\] \[y_2=0+0.4f(1,0)\] \[y_3=0+0.6f(1,0)\] \[y_4=0+0.8f(1,0)\] \[y_5=0+0.10f(1,0)\] \[y_6=0+0.12f(1,0)\] \[y_7=0+0.14f(1,0)\] Is this the idea?

OpenStudy (anonymous):

Ooops....It's supposed to + y_2 and y_3 and so on and so forth instead of 0's

OpenStudy (anonymous):

y_7=y_6+0.14f(1,0)

OpenStudy (anonymous):

oh no actually \( x_{n+1}=x_{n}+h \)

OpenStudy (anonymous):

I wrote the question wrong on here...it's supposed to be y'=y+xy

OpenStudy (anonymous):

no problem with that u have to find y(1.4) 1.4=1.2+0.2 =(1+0.2)+0.2 and note that actually \( y_{n}=y(x_{n}) \)